OpenStudy (anonymous):

Answer the questions about arc length If the circumference of the circle below is 25cm, what is the length of Arc AB, to the nearest hundredth?

OpenStudy (anonymous):



OpenStudy (nikato):

A circle is 360 And 70/360 = 7/36 So that "slice" is 7/36 of a circle So What's (7/36)(25)?

OpenStudy (caozeyuan):

so 2pir=25

OpenStudy (caozeyuan):

r=25/2pi

OpenStudy (caozeyuan):

s=r*theta

OpenStudy (caozeyuan):

so you can convert degrees to radians then you re done

OpenStudy (anonymous):

\[7\div36(25) equals 4.8611111111\] @nikato is this correct

OpenStudy (nikato):

Yes. That's what I got too

OpenStudy (anonymous):

ok thank you @nikato

OpenStudy (nikato):

No problem